OpenStudy (anonymous):

A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 25 feet and the child ride has a radius of 10 feet, what is the ratio between the volumes of each ride?

OpenStudy (dumbcow):

Volume of cone = 1/3pi*r^2*h height can be treated as function of radius so volume is a factor of radius cubed therefore ratio of volumes is ratio of radii cubed --> 10/25 = 2/5, (2/5)^3 = 8/125

OpenStudy (mertsj):

\[\frac{V _{1}}{V _{2}}=\frac{\pi25^2h}{\pi10^2h}=\frac{625}{100}=\frac{25}{4}\]

OpenStudy (anonymous):

Thanks, guys. :)

OpenStudy (mertsj):

yw